# Smishing Dataset I: Phishing SMS Dataset from Smishtank.com Data/Toolset paper

### v1

@@
Abstract While smishing (SMS Phishing) attacks have risen to become one of the most common types of social engineering attacks, there is a lack of relevant smishing datasets. One of the biggest challenges in the domain of smishing prevention is the availability of fresh smishing datasets. Additionally, as time persists, smishing campaigns are shut down and the crucial information related to the attack are lost. With the changing nature of smishing attacks, a consistent flow of new smishing examples is needed by both researchers and engineers to create effective defenses. In this paper, we present the community-sourced smishing datasets from the smishtank.com. It provides a wealth of information relevant to combating smishing attacks through the breakdown and analysis of smishing samples at the point of submission. In the contribution of our work, we provide a corpus of 1090 smishing samples that have been publicly submitted through the site. Each message includes information relating to the sender, message body, and any brands referenced in the message. Additionally, when a URL is found, we provide additional information on the domain, VirusTotal results, and a characterization of the URL. Through the open access of fresh smishing data, we empower academia and industries to create robust defenses against this evolving threat.
@@ 1. Introduction
Smishing, a variant of phishing that leverages text messages, has experienced a notable surge in frequency over recent years (FCCRobotext, ). Responding to this escalating threat, the Federal Communications Commission (FCC) took an unprecedented measure by instructing mobile providers to proactively block messages that are deemed ”highly likely to be illegal. Prior research on smishing attacks indicates that victims often exhibit a high susceptibility, frequently engaging with these campaigns by clicking on links, responding to the texts, or calling numbers included in the messages (Raman2023, ; timkounveil23, ). What makes phishing, and in particular smishing, attacks so difficult to prevent is their ever changing nature. Year over year, smishing trends and strategies change, introducing novel smishing attacks. Examples of this trend include a significant rise in COVID-related smishing incidents during the peak periods of the pandemic (irswarn, ).
@@
As part of this approach, we present a dataset comprising 1090 phishing SMSes sourced from smishtank.com. smishtank.com aims to address the scarcity of relevant public data by continually updating fresh SMS phishing messages. Leveraging open-source and public API resources, Smishtank consolidates varied sources into a single accessible platform. Presently, no other websites offer users the opportunity to publicly submit smishing messages for collaborative dataset creation.
@@
    We make publicly available a corpus of over 1090 smishing messages.
@@ 3.1. Message Collection
Messages were gathered exclusively from the smishing submission section on smishtank.com. These submissions comprise either a screenshot of the message or its text copy. Community-sourced datasets have become a prevalent and valuable resource in phishing research. In the realm of phishing URLs, Phishtank.com (phishtank, ) stands as an essential repository for new phishing URLs contributed by users. To augment submissions on the site, we promoted its features across social media platforms, encouraging users to share instances of smishing messages they received. Currently, we’ve amassed a total of 1090 validated submissions since the site’s launch. Upon submission, these screenshots are stored in a data lake in a standard image format, with the extracted data incorporated into the dataset.
@@ 4. Message Analysis
Figure 3. 10 most frequent imitated brands in dataset.
@@ 4.5. Top Level Domains
The TLD holds significant importance as an indicator for a suspicious website (Christou_2020, ). These domains represent the highest level in the domain hierarchy and appear following the final dot in a website’s domain. Phishing domains frequently imitate authentic ones by employing similar names with different TLDs. In Figure 6, we showcase a list of the ten most commonly encountered TLDs within our dataset.
Figure 6. Top Level Domains
@@ 4.6. VirusTotal Scores
Scores >=1 >=5 >=10 >=15
Detected 484 203 125 93
Phishing 368 158 95 23
Table 1. VirusTotal Scores
Figure 7. Domain Registrars
@@ 7. Conclusion
As smishing attacks evolve, the research community requires an increasing number of smishing data to effectively counter these social engineering attacks. Our efforts have involved collecting, processing, and categorizing 1090 phishing messages, aiming to empower developers and researchers in enhancing detection methods and advancing research on smishing campaigns. To ensure the timeliness of these messages, we’ve incorporated WHOIS and VirusTotal information, offering insights into their history at the point of collection. We anticipate that this work will lay a crucial foundation for future smishing research.
@@ 8. Acknowledgement
The authors thank all participants who submitted smishing samples to https://smishtank.com, thus enhancing our smishing dataset.

### v2

@@
Abstract.
While smishing (SMS Phishing) attacks have risen to become one of the most common types of social engineering attacks, there is a lack of relevant smishing datasets. One of the biggest challenges in the domain of smishing prevention is the availability of fresh smishing datasets. Additionally, as time persists, smishing campaigns are shut down and the crucial information related to the attack are lost. With the changing nature of smishing attacks, a consistent flow of new smishing examples is needed by both researchers and engineers to create effective defenses. In this paper, we present the community-sourced smishing datasets from the smishtank.com. It provides a wealth of information relevant to combating smishing attacks through the breakdown and analysis of smishing samples at the point of submission. In the contribution of our work, we provide a corpus of 1062 smishing samples that have been publicly submitted through the site. Each message includes information relating to the sender, message body, and any brands referenced in the message. Additionally, when a URL is found, we provide additional information on the domain, VirusTotal results, and a characterization of the URL. Through the open access of fresh smishing data, we empower academia and industries to create robust defenses against this evolving threat.
Smishing, Dataset, Phishing, VirusTotal
journalyear: 2024copyright: acmlicensedconference: Proceedings of the Fourteenth ACM Conference on Data and Application Security and Privacy; June 19–21, 2024; Porto, Portugalbooktitle: Proceedings of the Fourteenth ACM Conference on Data and Application Security and Privacy (CODASPY ’24), June 19–21, 2024, Porto, Portugaldoi: 10.1145/3626232.3653282isbn: 979-8-4007-0421-5/24/06ccs: Security and privacy Usability in security and privacyccs: Security and privacy Phishingccs: Security and privacy Social aspects of security and privacy
@@ 1. Introduction
Smishing, a variant of phishing that leverages text messages, has experienced a notable surge in frequency over recent years (FCCRobotext, ). Responding to this escalating threat, the Federal Communications Commission (FCC) took an unprecedented measure by instructing mobile providers to proactively block messages that are deemed highly likely to be illegal (FCCRulesChange, ). Prior research on smishing attacks indicates that victims often exhibit a high susceptibility, frequently engaging with these campaigns by clicking on links, responding to the texts, or calling numbers included in the messages (Raman2023, ; timkounveil23, ). What makes phishing, and in particular smishing, attacks so difficult to prevent is their ever changing nature. Year over year, smishing trends and strategies change, introducing novel smishing attacks. Examples of this trend include a significant rise in COVID-related smishing incidents during the peak periods of the pandemic (irswarn, ).
@@
As part of this approach, we present a dataset comprising 1062 phishing SMSes sourced from smishtank.com. smishtank.com aims to address the scarcity of relevant public data by continually updating fresh SMS phishing messages. Leveraging open-source and public API resources, Smishtank consolidates varied sources into a single accessible platform. Presently, no other websites offer users the opportunity to publicly submit smishing messages for collaborative dataset creation.
@@
    We make publicly available a corpus of 1062 smishing messages.
@@ 3.1. Message Collection
Messages were gathered exclusively from the smishing submission section on smishtank.com. These submissions comprise either a screenshot of the message or its text copy. Community-sourced datasets have become a prevalent and valuable resource in phishing research. In the realm of phishing URLs, Phishtank.com (phishtank, ) stands as an essential repository for new phishing URLs contributed by users. To augment submissions on the site, we promoted its features across social media platforms, encouraging users to share instances of smishing messages they received. Currently, we’ve amassed a total of 1062 validated submissions since the site’s launch. Upon submission, these screenshots are stored in a data lake in a standard image format, with the extracted data incorporated into the dataset.
@@ 4. Message Analysis
Figure 2. SMS Sender proportions by type.
@@ 4.1. Sender Information in Messages
Figure 3. 10 most frequent imitated brands in dataset.
@@ 4.5. Top Level Domains
The TLD holds significant importance as an indicator for a suspicious website (Christou_2020). These domains represent the highest level in the domain hierarchy and appear following the final dot in a website’s domain. Phishing domains frequently imitate authentic ones by employing similar names with different TLDs. In Figure 6, we showcase a list of the ten most commonly encountered TLDs within our dataset.
Scores >=1 >=5 >=10 >=15
Detected 486 203 125 93
Phishing 371 158 95 23
Table 1. VirusTotal Scores
@@ 7. Conclusion
As smishing attacks evolve, the research community requires an increasing number of smishing data to effectively counter these social engineering attacks. Our efforts have involved collecting, processing, and categorizing 1062 phishing messages, aiming to empower developers and researchers in enhancing detection methods and advancing research on smishing campaigns. To ensure the timeliness of these messages, we’ve incorporated WHOIS and VirusTotal information, offering insights into their history at the point of collection. We anticipate that this work will lay a crucial foundation for future smishing research.
@@ 8. Acknowledgement
The authors thank all participants who submitted smishing samples to https://smishtank.com, thus enhancing our smishing dataset. We thank all the CODASPY2̀4 reviewers for their thoughtful feedback and suggestions.